40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4047     Entry time: Mon Dec 13 18:06:43 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMY resuspended, ready to install. Tip Tilt realigned, ready to install. 

[Koji, Jenne]

I wish I could use a bigger font for this, but, the suspension work is totally done for the upgrade!!!

Now, nobody break any suspensions, or we're not going to be friends for a while. 

Koji and I put ETMY back in its tower, and made sure that both scribe lines are at the correct height.  We also confirmed that the balance is good (as Suresh mentioned in a previous elog, since we balanced using the AR surface, the HR surface is pointing downward a little bit, but it's well within the OSEMs ability to correct.

While we were in there, we also looked at Tip Tilt number 002.  As mentioned in elog 3645, the pitch pointing was off by a little bit.  Since the TTs don't have actuators, the pointing has to be pretty good.  We tweaked the balancing, and now the reflected beam goes completely back into the laser aperture, so it's as balanced as it's going to get.  This TT is now ready for installation onto the ITMY table as part of the SRC.

Kiwamu confirmed that he's going to install these optics tomorrow, since he's doing some other alignment work today.

Just for good measure, the Table:

StatusTable.png

ELOG V3.1.3-